I have to solve the problem: 5/8x + 3/5 = 8
I know I have to multiply each side by 5 but I'm lost after that. Any help would be great!

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Next step: Multiply both sides by


John


Answer by stanbon(75887)

I have to solve the problem: 5/8x + 3/5 = 8
I know I have to multiply each side by 5 but I'm lost after that. Any help would be great!
------------------
(5/8)x + (3/5) = 8
The lcm (least common multiple) of the denominators as 40.
Multiply thru by 40 to get rid of the denominators:
25x + 24 = 320
25x = 296
x = 296/25
======================
Cheers,
Stan H.